# Logarithmic Regret for Online KL-Regularized Reinforcement Learning

### v3

@@ 1 Introduction
We study the KL-regularized contextual bandit (Langford and Zhang, 2007; Xiong et al., 2024a) and Markov decision processes (MDPs) (Sutton, 2018) in this paper. These two frameworks have received significant attention due to their tremendous successes in the post-training stage of modern large language models (LLMs), which is commonly referred to as the Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022; Ouyang et al., 2022). The RLHF learning paradigm has been used to make a powerful chatbot by aligning the LLMs with human preference, making the model generation helpful, honest, and harmless (Bai et al., 2022). Notable examples include the OpenAI’s Chat-GPT (OpenAI, 2023), Anthropic’s Claude (Bai et al., 2022), and Google’s Gemini (Team et al., 2023). More recently, the KL-regularized RL framework has also been applied to enhance multi-turn reasoning capabilities, resulting in powerful reasoning models such as GPT4-o1 and DeepSeek-R1.
@@
However, despite the empirical success, a fundamental connection between KL regularization and learning efficiency has not been established so far theoretically. While some recent studies have attempted to analyze these frameworks, they often rely on standard analysis techniques, yielding regret guarantees similar to those of standard RL (Xiong et al., 2024a; Xie et al., 2024; Xiong et al., 2024b; Cen et al., 2024), or they depend on strong coverage assumptions and restricted to the bandit settings (Zhao et al., 2024). The following pivotal question remains open:
@@
Bandits Online Iterative GSHF [^1] Xiong et al. (2024a) studies relative-preference feedback, so their sample complexity additionally depends on $e^{\eta}$. Since our work focuses on absolute rewards, we omit this dependence. (Xiong et al., 2024a) -- $\widetilde{\mathcal{O}}\big{(}d^{2}/{\color[rgb]{1,0,0}\epsilon^{2}}\big{)}$$\times$
Two-Stage Mixed-Policy Sampling (Zhao et al., 2024) -- $\widetilde{\mathcal{O}}\big{(}(\eta^{2}C_{\mathrm{GL}}^{2}+\eta/{\color[rgb]{1,0,0}\epsilon})\log N_{\mathcal{R}}\big{)}$
KL-UCB (Theorem 4.1) $\mathcal{O}\big{(}\eta d_{\mathcal{R}}{\color[rgb]{1,0,0}\log T}\cdot\log(N_{\mathcal{R}})\big{)}$ $\widetilde{\mathcal{O}}\big{(}\eta d_{\mathcal{R}}\cdot\log(N_{\mathcal{R}})/{\color[rgb]{1,0,0}\epsilon}\big{)}$$\times$
MDPs Online Iterative M-GSHF [^2] Xiong et al. (2024b) considers the trajectory-level reward and learns the reward in $\mathcal{R}$ and transition probability in $\mathcal{P}$ separately, so the dependence on $d_{\mathcal{R}},N_{\mathcal{R}}$ and $d_{c}P,N_{\mathcal{P}}$ are separete. (Xiong et al., 2024b) $\widetilde{\mathcal{O}}\big{(}\sqrt{d_{\mathcal{R}}{\color[rgb]{1,0,0}T}\log N_{\mathcal{R}}}$ $+Hd_{\mathcal{P}}\log N_{\mathcal{P}}\big{)}$ $\widetilde{\mathcal{O}}\big{(}d_{\mathcal{R}}\log N_{\mathcal{R}}/{\color[rgb]{1,0,0}\epsilon^{2}}$ $+Hd_{\mathcal{P}}\log N_{\mathcal{P}}/\epsilon\big{)}$$\times$
XPO (Xie et al., 2024) [^3] Xie et al. (2024) considers the deterministic-transition setting and optimizes the policy directly. -- $\widetilde{\mathcal{O}}\big{(}d_{\mathcal{F}}\log N_{\mathcal{F}}/{\color[rgb]{1,0,0}\epsilon^{2}}\big{)}$$\times$
KL-LSVI-UCB (Theorem 5.1) $\mathcal{O}\big{(}\eta H^{2}d_{\mathcal{F}}{\color[rgb]{1,0,0}\log T}\cdot\log N_{\mathcal{F}\oplus\mathcal{B}})\big{)}$ $\widetilde{\mathcal{O}}\big{(}\eta H^{2}d_{\mathcal{F}}\cdot\log N_{\mathcal{F}\oplus\mathcal{B}})/{\color[rgb]{1,0,0}\epsilon}\big{)}$$\times$
2 Related Work
@@
The theoretical foundation of RLHF traces back to dueling bandits (e.g., Yue et al., 2012; Saha, 2021; Bengs et al., 2021), which studied preference feedback in non-regularized settings. This was later extended to online reinforcement learning with finite state spaces (tabular settings) and function approximation (Xu et al., 2020; Novoseller et al., 2020; Pacchiano et al., 2021; Chen et al., 2022). More recently, Zhan et al. (2023b); Wu and Sun (2023) developed reward-free learning algorithms and sampling-based methods for online RLHF. For the offline learning, Zhu et al. (2023); Zhan et al. (2023a); Li et al. (2023b); Zhong et al. (2024); Huang et al. (2024) propose sample-efficient algorithms under suitable coverage conditions. These works mainly develop techniques to estimate the underlying reward model associated with the Bradley-Terry model from querying the preference oracle (human) and achieve similar order regret with the standard reward learning. However, since they only consider reward maximization, the results deviate from the practical applications of RLHF. For example, reward maximization frameworks often assume a deterministic optimal policy, which is unsuitable for generative models.
@@ 3.1 KL-Regularized Contextual Bandits
$U_{\mathcal{R}}(\lambda,x,a;\mathcal{D}_{t-1})=\sup_{R_{1},R_{2}\in\mathcal{R}}\frac{|R_{1}(x,a)-R_{2}(x,a)|}{\sqrt{\lambda+\sum_{i=1}^{t-1}\big{(}R_{1}(x_{i},a_{i})-R_{2}(x_{i},a_{i})\big{)}^{2}}}.$
@@
where the inequality uses the Cauchy–Schwarz inequality. Hence, the uncertainty reflects how much a direction in the feature space has been explored. Further, the eluder dimension represents how many times the model can be “surprised” by the out-of-sample data over $T$ rounds. One can show that this definition is more general than the original one in Russo and Van Roy (2013) as we further take the magnitude of violation into consideration while the original eluder dimension only counts the frequency (Zhang, 2023; Xie et al., 2022).
@@
Lemma 3.4 (Solution of KL-regularized Optimization (Proposition 7.16 and Theorem 15.3 of Zhang 2023)).
@@
$\max_{\pi}\mathbb{E}_{a\sim\pi(\cdot|x)}\Big{[}R(x,a)-\eta^{-1}\mathrm{KL}\big{(}\pi(\cdot|x)\|\pi_{\mathrm{ref}}(\cdot|x\big{)}\Big{]}=\frac{1}{\eta}\cdot\log\mathbb{E}_{a\sim\pi_{\mathrm{ref}}(\cdot|x)}\exp\big{(}\eta R(x,a)\big{)},$
@@ 3.2 KL-Regularized Reinforcement Learning
$V^{\pi}_{h}(s_{h})=$ $\sum_{h^{\prime}=h}^{H}\mathbb{E}^{\pi}\Big{[}r_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})-\frac{1}{\eta}\cdot\mathrm{KL}\bigl{(}\pi_{h^{\prime}}(\cdot|s_{h^{\prime}})\|\pi_{\mathrm{ref},h^{\prime}}(\cdot|s_{h^{\prime}})\bigr{)}\,\Big{|}\,s_{h}\Big{]},$
$Q^{\pi}_{h}(s_{h},a_{h})=$ $r_{h}(s_{h},a_{h})+\mathbb{E}^{\pi}\sum_{h^{\prime}=h+1}^{H}\Big{[}r_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})-\frac{1}{\eta}\cdot\mathrm{KL}\bigl{(}\pi_{h^{\prime}}(\cdot|s_{h^{\prime}})\|\pi_{\mathrm{ref},h^{\prime}}(\cdot|s_{h^{\prime}})\bigr{)}\,\Big{|}\,s_{h},a_{h}\Big{]}.$
@@
$V_{h}^{\pi}(s_{h})=\mathbb{E}_{a_{h}\sim\pi_{h}(\cdot|s_{h})}\Big{[}Q_{h}^{\pi}(s_{h},a_{h})-\frac{1}{\eta}\cdot\mathrm{KL}\bigl{(}\pi_{h}(\cdot|s_{h})\|\pi_{\mathrm{ref},h}(\cdot|s_{h})\bigr{)}\Big{]},$
@@
$V_{h}^{*}(s_{h})=\frac{1}{\eta}\log\mathbb{E}_{a_{h}\sim\pi_{\mathrm{ref},h}(\cdot|s_{h})}\exp\big{(}\eta Q_{h}^{*}(s_{h},a_{h})\big{)}.$ (3.1)
@@
$Q_{h}^{*}(s_{h},a_{h})=r_{h}(s_{h},a_{h})+\mathbb{E}_{s_{h+1}\sim\mathbb{P}_{h}(\cdot|s_{h},a_{h})}V_{h+1}^{*}(s_{h+1}).$ (3.2)
Without loss of generality, we assume that $Q_{h}^{*}(s_{h},a_{h})\in[0,1]$ for any $(s_{h},a_{h})\in{\mathcal{S}}\times\mathcal{A}$ [^1] This is for the simplicity of analysis. By multiplying the results by $H$, our analysis can be applied to the cases where $r_{h}\in[0,1]$ for $h\in[H]$.. From (3.1), we know that $V_{h}^{*}(s_{h})\in[0,1]$ for any $s_{h}\in\mathcal{A}$.
@@ 4.1 The Proposed Algorithm and Result
6:     Apply the planning oracle to compute $\pi_{t+1}(\cdot|\cdot)\propto\pi_{\mathrm{ref}}(\cdot|\cdot)\exp\bigl{[}\eta\big{(}\widehat{R}_{t}(\cdot,\cdot)+b_{t}(\cdot,\cdot)\big{)}\bigr{]}$, where $b_{t}$ is the exploration bonus term defined in (4.1).
@@
$b_{t}(x,a)=\min\Big{\{}1,\beta_{T}\cdot U_{\mathcal{R}_{t}}(\lambda,x,a;\mathcal{D}_{t})\Big{\}},$ (4.1)
@@
Review of Analysis in Previous Work.
@@
$=\underbrace{R(x,\pi^{*})-\widehat{R}_{t}(x,\pi^{*})-b_{t}(x,\pi^{*})}_{\displaystyle{\leq 0~{}\text{(By Optimism)}}}+\underbrace{\widehat{R}_{t}(x,\pi_{t})-R^{*}(x,\pi_{t})+b_{t}(x,\pi_{t})}_{\displaystyle{\leq 2b_{t}(x,\pi_{t})}}$
@@
$\mathbb{E}_{\pi^{*}}\bigg{[}R^{*}(x,a)-\frac{1}{\eta}\log\frac{\pi^{*}(a|x)}{\pi_{\mathrm{ref}}(a|x)}\bigg{]}-\mathbb{E}_{\pi_{t}}\bigg{[}R^{*}(x,a)-\frac{1}{\eta}\log\frac{\pi_{t}(a|x)}{\pi_{\mathrm{ref}}(a|x)}\bigg{]}$ (4.2)
$=\frac{1}{\eta}\log Z_{R^{*}}(x)-\frac{1}{\eta}\log Z_{\widehat{R}_{t-1}+b_{t-1}}(x)+\mathbb{E}_{\pi_{t}}\big{[}\widehat{R}_{t-1}(x,a)-R^{*}(x,a)+b_{t-1}(x,a)\big{]},$
where $Z_{R}(x)$ is the normalization term for the policy with respect to reward function $R$. We defer the detailed derivation of (4.2) to Appendix A.
@@
Compared to previous work (Xiong et al., 2024a; Xie et al., 2024; Ye et al., 2024) that neglect the KL term and reduce to the traditional RL analysis, the upper bound in (4.2) is more refined and involves the gap of the normalization term $Z_{R}(x)$, which is crucial for the analysis of the KL-regularized objective.
@@
To further analyze (4.2), we define a function
$\Delta(x,R):=-\frac{1}{\eta}\log Z_{R}(x)+\mathbb{E}_{\pi_{R}^{\eta}}\big{[}R(x,a)-R^{*}(x,a)\big{]}.$ (4.3)
and write (4.2) as the function gap:
$\mathbb{E}_{x\sim d_{0}}[\Delta(x,\widehat{R}_{t-1}+b_{t-1})-\Delta(x,R^{*})].$ (4.4)
To analyze the behavior of (4.4), we have the following lemma, which reveals the gradient of $\Delta(x,R)$ with respect to $R$.
@@
$\frac{\partial\Delta(x,R)}{\partial R(x,a)}=\eta\pi_{R}^{\eta}(a|x)\cdot\big{(}R(x,a)-R^{*}(x,a)\big{)}$ (4.5)
@@
$\mathbb{E}_{x\sim d_{0}}[\Delta(x,\widehat{R}_{t-1}+b_{t-1})-\Delta(x,R^{*})]=\mathbb{E}_{x\sim d_{0}}\Big{[}\sum_{a\in\mathcal{A}}\Big{(}\eta\pi_{f_{\lambda}}^{\eta}(a|x)\cdot\big{(}f_{\lambda}(x,a)-R^{*}(x,a)\big{)}$
$\qquad-\eta\sum_{a^{\prime}\in\mathcal{A}}\pi_{f_{\lambda}}^{\eta}(a|x)\cdot\pi_{f_{\lambda}}^{\eta}(a^{\prime}|x)\cdot\big{(}f_{\lambda}(x,a^{\prime})-R^{*}(x,a^{\prime})\big{)}\Big{)}\cdot\big{(}\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)-R^{*}(x,a)\big{)}\Big{]},$
@@
$\mathbb{E}_{x\sim d_{0}}[\Delta(x,\widehat{R}_{t-1}+b_{t-1})-\Delta(x,R^{*})]$ (4.6)
$\leq\eta\mathbb{E}_{x\sim d_{0}}\big{[}\sum_{a\in\mathcal{A}}\bigl{(}\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)-R^{*}(x,a)\bigr{)}\cdot\underbrace{\bigl{(}f_{\lambda}(x,a)-R^{*}(x,a)\bigr{)}\cdot\pi_{f_{\lambda}}^{\eta}(a|x)}_{U(x,a)}\big{]}$
$\leq\eta\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{\widehat{R}_{t-1}+b_{t-1}}^{\eta}(\cdot|x)}\big{[}\bigl{(}\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)-R^{*}(x,a)\bigr{)}^{2}\big{]},$
@@
Conditioning on high-probability event $\cup_{t\in[T]}\mathcal{E}_{\mathrm{opt}}^{t}$, we derive the regret bound from (4.6):
$\mathrm{Regret}(T)\leq\eta\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim d_{0}}\mathbb{E}_{a_{t}\sim\pi_{t}}\big{[}\bigl{(}\widehat{R}_{t-1}(x_{t},a_{t})+b_{t-1}(x_{t},a_{t})-R^{*}(x_{t},a_{t})\bigr{)}^{2}\big{]}$
@@
Thus, according to the definition of bonus $b_{t}$ in (4.1), we can follow standard analyses for general function approximation under the finite eluder dimension to obtain the desired result.

### v5

@@ 1 Introduction
We study the KL-regularized contextual bandit (Langford & Zhang, 2007; Xiong et al., 2024a) and Markov decision processes (MDPs) (Sutton, 2018) in this paper. These two frameworks have received significant attention due to their tremendous successes in the post-training stage of modern large language models (LLMs), which is commonly referred to as the Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022; Ouyang et al., 2022). The RLHF learning paradigm has been used to make a powerful chatbot by aligning the LLMs with human preference, making the model generation helpful, honest, and harmless (Bai et al., 2022). Notable examples include the OpenAI’s Chat-GPT (OpenAI, 2023), Anthropic’s Claude (Bai et al., 2022), and Google’s Gemini (Team et al., 2023). More recently, the KL-regularized RL framework has also been applied to enhance multi-turn reasoning capabilities, resulting in powerful reasoning models such as GPT4-o1 and DeepSeek-R1.
@@
However, despite the empirical success, a fundamental connection between KL regularization and learning efficiency has not been established so far theoretically. While some recent studies have attempted to analyze these frameworks, they often rely on standard analysis techniques, yielding regret guarantees similar to those of standard RL (Xiong et al., 2024a; Xie et al., 2024; Xiong et al., 2024b; Cen et al., 2024), or they depend on strong coverage assumptions and restricted to the bandit settings (Zhao et al., 2024). A notable exception is the approach of Tiapkin et al. (2023, 2024), which achieves $\widetilde{\mathcal{O}}(H^{5}d^{2}\eta/\epsilon)$ sample complexity for KL-regularized linear MDPs, and $\widetilde{\mathcal{O}}(H^{5}S^{2}A\eta/\epsilon)$ sample complexity for tabular MDPs. However, their analysis focuses on the pure exploration or best policy identification setting, where the goal is to find near-optimal policies using the least possible amount of interactions with the environment, rather than the online setting where the agent needs to trade off between exploration and exploitation. Therefore, the following pivotal question remains open:
@@ 2 Related Work
Bandits Online Iterative GSHF [^1] Xiong et al. (2024a) studies relative-preference feedback, so their sample complexity additionally depends on $e^{\eta}$. Since our work focuses on absolute rewards, we omit this dependence. (Xiong et al., 2024a) -- $\widetilde{\mathcal{O}}\big{(}d^{2}/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon^{2}}\big{)}$$\times$
Two-Stage Mixed-Policy Sampling (Zhao et al., 2024) -- $\widetilde{\mathcal{O}}\big{(}(\eta^{2}C_{\mathrm{GL}}^{2}+\eta/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon})\log N_{\mathcal{R}}\big{)}$
KL-UCB (Theorem 4.1) $\mathcal{O}\big{(}\eta d_{\mathcal{R}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\log T}\cdot\log(N_{\mathcal{R}})\big{)}$ $\widetilde{\mathcal{O}}\big{(}\eta d_{\mathcal{R}}\cdot\log(N_{\mathcal{R}})/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon}\big{)}$$\times$
MDPs Online Iterative M-GSHF [^2] Xiong et al. (2024b) considers the trajectory-level reward and learns the reward in $\mathcal{R}$ and transition probability in $\mathcal{P}$ separately, so the dependence on $d_{\mathcal{R}},N_{\mathcal{R}}$ and $d_{c}P,N_{\mathcal{P}}$ are separete. (Xiong et al., 2024b) $\widetilde{\mathcal{O}}\big{(}\sqrt{d_{\mathcal{R}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}T}\log N_{\mathcal{R}}}$ $+Hd_{\mathcal{P}}\log N_{\mathcal{P}}\big{)}$ $\widetilde{\mathcal{O}}\big{(}d_{\mathcal{R}}\log N_{\mathcal{R}}/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon^{2}}$ $+Hd_{\mathcal{P}}\log N_{\mathcal{P}}/\epsilon\big{)}$$\times$
XPO (Xie et al., 2024) [^3] Xie et al. (2024) considers the deterministic-transition setting and optimizes the policy directly. -- $\widetilde{\mathcal{O}}\big{(}d_{\mathcal{F}}\log N_{\mathcal{F}}/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon^{2}}\big{)}$$\times$
KL-LSVI-UCB (Theorem 5.1) $\mathcal{O}\big{(}\eta H^{2}d_{\mathcal{F}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\log T}\cdot\log N_{\mathcal{F}\oplus\mathcal{B}})\big{)}$ $\widetilde{\mathcal{O}}\big{(}\eta H^{2}d_{\mathcal{F}}\cdot\log N_{\mathcal{F}\oplus\mathcal{B}})/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon}\big{)}$$\times$
@@
The theoretical foundation of RLHF traces back to dueling bandits (e.g., Yue et al., 2012; Saha, 2021; Bengs et al., 2021), which studied preference feedback in non-regularized settings. This was later extended to online reinforcement learning with finite state spaces (tabular settings) and function approximation (Xu et al., 2020; Novoseller et al., 2020; Pacchiano et al., 2021; Chen et al., 2022). More recently, Zhan et al. (2023b); Wu & Sun (2023) developed reward-free learning algorithms and sampling-based methods for online RLHF. For the offline learning, Zhu et al. (2023); Zhan et al. (2023a); Li et al. (2023b); Zhong et al. (2024); Huang et al. (2024) propose sample-efficient algorithms under suitable coverage conditions. These works mainly develop techniques to estimate the underlying reward model associated with the Bradley-Terry model from querying the preference oracle (human) and achieve similar order regret with the standard reward learning. However, since they only consider reward maximization, the results deviate from the practical applications of RLHF. For example, reward maximization frameworks often assume a deterministic optimal policy, which is unsuitable for generative models.
@@ 3.1 KL-Regularized Contextual Bandits
$U_{\mathcal{R}}(\lambda,x,a;\mathcal{D}_{t-1})$
$=\sup_{R_{1},R_{2}\in\mathcal{R}}\frac{|R_{1}(x,a)-R_{2}(x,a)|}{\sqrt{\lambda+\sum_{i=1}^{t-1}\big{(}R_{1}(x_{i},a_{i})-R_{2}(x_{i},a_{i})\big{)}^{2}}}.$
@@
where the inequality uses the Cauchy–Schwarz inequality. Hence, the uncertainty reflects how much a direction in the feature space has been explored. Further, the eluder dimension represents how many times the model can be “surprised” by the out-of-sample data over $T$ rounds. One can show that this definition is more general than the original one in (Russo & Van Roy, 2013) as we further take the magnitude of violation into consideration while the original eluder dimension only counts the frequency (Zhang, 2023; Xie et al., 2022).
@@
Lemma 3.4 (Solution of KL-regularized Optimization (Proposition 7.16 and Theorem 15.3 of Zhang (2023)).
@@
$\max_{\pi}\mathbb{E}_{a\sim\pi(\cdot|x)}\Big{[}R(x,a)-\eta^{-1}\mathrm{KL}\big{(}\pi(\cdot|x)\|\pi_{\mathrm{ref}}(\cdot|x\big{)}\Big{]}$
$=\frac{1}{\eta}\cdot\log\mathbb{E}_{a\sim\pi_{\mathrm{ref}}(\cdot|x)}\exp\big{(}\eta R(x,a)\big{)},$
@@ 3.2 KL-Regularized Reinforcement Learning
$V^{\pi}_{h}(s_{h})=$ $\sum_{h^{\prime}=h}^{H}\mathbb{E}^{\pi}\Big{[}r_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})$
$\qquad-\frac{1}{\eta}\cdot\mathrm{KL}\bigl{(}\pi_{h^{\prime}}(\cdot|s_{h^{\prime}})\|\pi_{\mathrm{ref},h^{\prime}}(\cdot|s_{h^{\prime}})\bigr{)}\,\Big{|}\,s_{h}\Big{]},$
$Q^{\pi}_{h}(s_{h},a_{h})=$ $r_{h}(s_{h},a_{h})+\mathbb{E}^{\pi}\sum_{h^{\prime}=h+1}^{H}\Big{[}r_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})$
$\qquad-\frac{1}{\eta}\cdot\mathrm{KL}\bigl{(}\pi_{h^{\prime}}(\cdot|s_{h^{\prime}})\|\pi_{\mathrm{ref},h^{\prime}}(\cdot|s_{h^{\prime}})\bigr{)}\,\Big{|}\,s_{h},a_{h}\Big{]}.$
@@
$V_{h}^{\pi}(s_{h})=\mathbb{E}_{a_{h}\sim\pi_{h}(\cdot|s_{h})}\Big{[}Q_{h}^{\pi}(s_{h},a_{h})$
$\qquad\qquad-\frac{1}{\eta}\cdot\mathrm{KL}\bigl{(}\pi_{h}(\cdot|s_{h})\|\pi_{\mathrm{ref},h}(\cdot|s_{h})\bigr{)}\Big{]},$
@@
$\small V_{h}^{*}(s_{h})=\frac{1}{\eta}\log\mathbb{E}_{a_{h}\sim\pi_{\mathrm{ref},h}(\cdot|s_{h})}\exp\big{(}\eta Q_{h}^{*}(s_{h},a_{h})\big{)}.$ (1)
@@
$\small Q_{h}^{*}(s_{h},a_{h})=r_{h}(s_{h},a_{h})+\mathbb{E}_{s_{h+1}\sim\mathbb{P}_{h}(\cdot|s_{h},a_{h})}V_{h+1}^{*}(s_{h+1}).$ (2)
Without loss of generality, we assume that $Q_{h}^{*}(s_{h},a_{h})\in[0,1]$ for any $(s_{h},a_{h})\in{\mathcal{S}}\times\mathcal{A}$ [^1] This is for the simplicity of analysis. By multiplying the results by $H$, our analysis can be applied to the cases where $r_{h}\in[0,1]$ for $h\in[H]$.. From (1), we know that $V_{h}^{*}(s_{h})\in[0,1]$ for any $s_{h}\in\mathcal{A}$.
@@ 4.1 The Proposed Algorithm and Result
6:     Apply the planning oracle to compute $\pi_{t+1}(\cdot|\cdot)\propto\pi_{\mathrm{ref}}(\cdot|\cdot)\exp\bigl{[}\eta\big{(}\widehat{R}_{t}(\cdot,\cdot)+b_{t}(\cdot,\cdot)\big{)}\bigr{]}$, where $b_{t}$ is the exploration bonus term defined in (3).
@@
$b_{t}(x,a)=\min\Big{\{}1,\beta_{T}\cdot U_{\mathcal{R}_{t}}(\lambda,x,a;\mathcal{D}_{t})\Big{\}},$ (3)
@@
Review of Analysis in Previous Work
@@
$=\underbrace{R(x,\pi^{*})-\widehat{R}_{t}(x,\pi^{*})-b_{t}(x,\pi^{*})}_{\displaystyle{\leq 0~{}\text{(By Optimism)}}}$
$\qquad+\underbrace{\widehat{R}_{t}(x,\pi_{t})-R^{*}(x,\pi_{t})+b_{t}(x,\pi_{t})}_{\displaystyle{\leq 2b_{t}(x,\pi_{t})}}$
@@
$\mathbb{E}_{\pi^{*}}\bigg{[}R^{*}(x,a)-\frac{1}{\eta}\log\frac{\pi^{*}(a|x)}{\pi_{\mathrm{ref}}(a|x)}\bigg{]}$ (4)
$\quad-\mathbb{E}_{\pi_{t}}\bigg{[}R^{*}(x,a)-\frac{1}{\eta}\log\frac{\pi_{t}(a|x)}{\pi_{\mathrm{ref}}(a|x)}\bigg{]}$
$=\frac{1}{\eta}\log Z_{R^{*}}(x)-\frac{1}{\eta}\log Z_{\widehat{R}_{t-1}+b_{t-1}}(x)$
$\quad+\mathbb{E}_{\pi_{t}}\big{[}\widehat{R}_{t-1}(x,a)-R^{*}(x,a)+b_{t-1}(x,a)\big{]},$
where $Z_{R}(x)$ is the normalization term for the policy with respect to reward function $R$. We defer the detailed derivation of (4) to Appendix A.
@@
Compared to previous work (Xiong et al., 2024a; Xie et al., 2024; Ye et al., 2024) that neglect the KL term and reduce to the traditional RL analysis, the upper bound in (4) is more refined and involves the gap of the normalization term $Z_{R}(x)$, which is crucial for the analysis of the KL-regularized objective.
@@
To further analyze (4), we define a function
$\Delta(x,R):=-\frac{1}{\eta}\log Z_{R}(x)+\mathbb{E}_{\pi_{R}^{\eta}}\big{[}R(x,a)-R^{*}(x,a)\big{]}.$ (5)
and write (4) as the function gap:
$\mathbb{E}_{x\sim d_{0}}[\Delta(x,\widehat{R}_{t-1}+b_{t-1})-\Delta(x,R^{*})].$ (6)
To analyze the behavior of (6), we have the following lemma, which reveals the gradient of $\Delta(x,R)$ with respect to $R$.
@@
$\frac{\partial\Delta(x,R)}{\partial R(x,a)}=\eta\pi_{R}^{\eta}(a|x)\cdot\big{(}R(x,a)-R^{*}(x,a)\big{)}$ (7)
@@
$=\mathbb{E}_{x\sim d_{0}}\Big{[}\sum_{a\in\mathcal{A}}\Big{(}\eta\pi_{f_{\lambda}}^{\eta}(a|x)\cdot\big{(}f_{\lambda}(x,a)-R^{*}(x,a)\big{)}$
$\qquad-\eta\sum_{a^{\prime}\in\mathcal{A}}\pi_{f_{\lambda}}^{\eta}(a|x)\cdot\pi_{f_{\lambda}}^{\eta}(a^{\prime}|x)\cdot\big{(}f_{\lambda}(x,a^{\prime})-R^{*}(x,a^{\prime})\big{)}\Big{)}$
$\qquad\qquad\cdot\big{(}\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)-R^{*}(x,a)\big{)}\Big{]},$
@@
$\mathbb{E}_{x\sim d_{0}}[\Delta(x,\widehat{R}_{t-1}+b_{t-1})-\Delta(x,R^{*})]$ (8)
$\leq\eta\mathbb{E}_{x\sim d_{0}}\big{[}\sum_{a\in\mathcal{A}}\bigl{(}\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)-R^{*}(x,a)\bigr{)}$
$\quad\cdot\underbrace{\bigl{(}f_{\lambda}(x,a)-R^{*}(x,a)\bigr{)}\cdot\pi_{f_{\lambda}}^{\eta}(a|x)}_{U(x,a)}\big{]}$
$\leq\eta\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{\widehat{R}_{t-1}+b_{t-1}}^{\eta}(\cdot|x)}\big{[}$
$\qquad\bigl{(}\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)-R^{*}(x,a)\bigr{)}^{2}\big{]},$
@@
Conditioning on high-probability event $\cup_{t\in[T]}\mathcal{E}_{\mathrm{opt}}^{t}$, we derive the regret bound from (8):
$\mathrm{Regret}(T)\leq\eta\sum_{t=1}^{T}\mathbb{E}_{x_{t}\sim d_{0}}\mathbb{E}_{a_{t}\sim\pi_{t}}\big{[}$
$\qquad\qquad\bigl{(}\widehat{R}_{t-1}(x_{t},a_{t})+b_{t-1}(x_{t},a_{t})-R^{*}(x_{t},a_{t})\bigr{)}^{2}\big{]}$
@@
Thus, according to the definition of bonus $b_{t}$ in (3), we can follow standard analyses for general function approximation under the finite eluder dimension to obtain the desired result.